Your company needs a new corporate jet for 5 years. You can either buy or lease it. If you lease it, you need to make a lease payment at the beginning of each year. All prices are in $ million.

A B
1 Price of jet 16
2 Resale value after 5 years 8
3 Annual lease payment 1.6
4 Number of lease payments 5
5 Interest rate 12%

Part 1

What is the net advantage to leasing, i.e., the incremental NPV of leasing relative to buying (in $ million)?
Answer #1

Buying -

Price of Jet now = CF0 = -16
Resale Value after 5 years = CF5 = 8
Interest rate = r = 12% or 0.12
NPV = CF0 + CF5/(1+r)5 = -16 + 8/1.125 = -11.46

Leasing -

Annual Lease Payment = 1.6
=> CF1 = CF2 = CF3 = CF4 = CF5 = -1.6
Interest Rate = r = 12% or 0.12
NPV = CF1/(1+r) + CF2/(1+r)2 + CF3/(1+r)3 + CF4/(1+r)4 + CF5/(1+r)5 = -1.6/1.12 - 1.6/1.122 - 1.6/1.123 - 1.6/1.124 - 1.6/1.125 = -5.77

Since the Present value of investment is lower for Leasing, leasing is a better option

Incremental NPV for leasing = 11.46 - 5.77 = $5.69m